What is the pH solution containing 0.30 M HCOOH and 0.52 M HCOOK?

HCOOH(aq) + H2O(l) = H3O+(aq) + HCOO-(aq)

1
Expert's answer
2021-05-03T07:25:53-0400

pKa = 3.77


pH = pKa + log ( "\frac{[HCOO^-]}{[HCOOH]}" )


pH = 3.77 + log ( "\frac{[0.52]}{[0.30]}" )


pH = 4.01